OpenStudy (anonymous):

To the nearest whole degree, what angle measure has a tangent of 9.1834?

OpenStudy (anonymous):

Maybe...83 degrees and 47 minutes = 9.1803

OpenStudy (jdoe0001):

\(\bf tan(\theta)=9.1834\implies tan^{-1}[tan(\theta)]=tan^{-1}(9.1834) \\ \quad \\ \theta=tan^{-1}(9.1834)\)